# Safe Planning in Unknown Environments using Conformalized Semantic Maps

### v1

@@ V-A Setting Up Experiments: Robot & Distribution $\mathcal{D}$}
Distribution $\mathcal{D}$}: The distribution $\mathcal{D}$} is designed to generate mission scenarios $\xi=\{{\mathbf{x}}_{0},\Omega,\psi\}$ as follows. All environments $\Omega$ have dimensions of $70\times 25$ m and contain objects from five semantic classes, $\mathcal{K}=\{\texttt{person},\texttt{car},\texttt{truck},\texttt{tree},\texttt{free}\}$. Specifically, each environment is populated with one car, one truck, one human, and three parallel rows of three trees. The tree locations remain fixed across all environments, while the car and truck are placed at randomized positions and orientations. The human position is also generated randomly, with a bias toward being placed near the trees. This choice enables testing our algorithm’s performance in cases of partial occlusions. The initial robot location ${\mathbf{x}}_{0}$ is selected randomly. The task $\psi$ is designed so that the goal location is chosen randomly, and the class–based safety distances $d_{k}$ in (9) are $d_{\text{Person}}=4$ m, $d_{\text{Car}}=1$ m, $d_{\text{Truck}}=2$ m, and $d_{\text{Tree}}=0.5$ m. An example mission scenario is shown in Figure 3.
V-B Setting Up the Proposed Planner & Evaluation Metrics
@@
Semantic Mapping: To set up the semantic mapping algorithm introduced in [19], we use a Detectron2 model [46], trained on the COCO dataset [47], to semantically segment RGB images. The segmented outputs are fused over time to construct a 3D semantic octomap with resolution $r_{m}=1$ m. In our experiments, since we focus on a ground robot, we project this 3D map onto the X–Y plane, while considering height ranges relevant to the robot and ignoring objects that do not interfere with its motion; this gives rise to a 2D semantic class–based grid map. In doing so, we assume that no semantic objects are stacked vertically.
@@
Exploitation-Exploration: The exploitation mode of Alg. 2 is implemented using a standard $\texttt{A}^{*}$ search algorithm on the 2D grid map [1]. For exploration, we design an algorithm that randomly selects a next state ${\mathbf{x}}_{t+1}$ near a voxel with a non-singleton prediction set subject to (6). This approach aims to decrease map uncertainty and specifically the size of prediction sets by driving the robot towards cell $j$ with non-singleton sets. As discussed in Section III-B, however, any other exploration algorithm can be substituted.
@@
Conformalization of Maps: Next, we discuss our implementation of the CP process described in Section III-A. Recall that the NCS in (4) requires evaluating the maximum over the set $\mathcal{P}$}, which contains an infinite number of paths. Since this is impractical, following [30], we approximate the NCS using a finite subset $\hat{\mathcal{P}}\subset{\mathcal{P}}$ consisting of 10 paths between the start and goal locations in a given calibration scenario $\xi_{i}$. Additionally, to avoid generating a new calibration dataset for every test scenario, we apply CP as described in Remark IV.4, with $\delta=0.1$. The value $\hat{\alpha}$ is computed such that $Beta_{D+1-v,v}^{-1}(\delta)\approx 1-\alpha$, and we then use $\hat{\alpha}$ to compute the quantile $\hat{s}$. Our calibration dataset consists of $50$ mission scenarios designed as discussed in Section V-A.
@@
Evaluation Metrics: We consider four metrics. (i) The first metric is the mapping success rate, defined as $S.R_{\text{map}}:=\mathds{1}(C_{0:H}),$ where $C_{0:H}=\bigwedge_{\forall j\in{\mathcal{J}}_{t},t\in[0:H]}\left(k_{j}^{\text{gt}}\in{\mathcal{C}}_{j}({\mathbf{m}}_{t})\right)$ and $k_{j}^{\text{gt}}$ is the ground truth label of cell $j$. In words, for a given test scenario $\xi_{\text{test}}$, $S.R_{\text{map}}=1$ if all constructed prediction sets ${\mathcal{C}}_{j}({\mathbf{m}}_{t})$ contain the true labels $k_{j}^{\text{gt}}$ for all $j\in{\mathcal{J}}_{t}$ and $t\in[0:H]$. Due to Proposition IV.1, the average mapping success rate of Alg. 2 across test scenarios drawn from $\mathcal{D}$} is expected to be at least $1-\alpha$. (ii) The second metric is the mission success rate, defined as $S.R_{\text{mission}}:=\mathds{1}\left({\mathbf{x}}_{0:H}\models\psi\right),$. In words, for a given test scenario $\xi_{\text{test}}$, $S.R_{\text{mission}}=1$ if the robot path ${\mathbf{x}}_{0:H}$ satisfies the safety constraints in (9) and ${\mathbf{x}}_{H}\in{\mathcal{X}}_{g}$. Due to Theorem IV.3, the average mission success rate of Alg. 2 across test scenarios drawn from $\mathcal{D}$} is expected to be at least $1-\alpha$. (iii) The third metric is the distance traveled by the robot during deployment, which captures efficiency of the generated paths. (iv) The fourth metric is the proportion of the overall path used for exploration, quantifying how much of the trajectory was dedicated to reducing uncertainty in the environment.
V-C Empirical Evaluation of the Proposed Planner
@@
We evaluate our proposed method for three user-defined success rates: $1-\alpha\in\{0.95,0.9,0.85\}$. Specifically, we consider $61$ test scenarios drawn from $\mathcal{D}$} and report the average mapping and mission success rates as defined above. Additionally, we report the average path length and the average proportion of the path used for exploration, computed over all successful mission scenarios. The empirical results are reported in Table I. Observe that the empirical mapping success rates for $1-\alpha=95\%,90\%$ and $85\%$ are $93.36\%,90.16\%$ and $86.89\%$ respectively, validating Proposition IV.1. Now, it can be seen that for $1-\alpha=0.95$, the empirical mapping success rate is $93.36\%<95\%$. Such small deviations are due to the finiteness of the calibration dataset, the number of paths that are considered for calibration, and the number of validation environments. For more information on the influence of finiteness on the empirical success rates, the readers are directed to [29]. In addition, the empirical mission success rates for $1-\alpha=95\%,90\%,85\%$ are $98.36\%,96.72\%,95.08\%$ respectively validating Theorem IV.3. Also, observe in Table I that the path length increases with higher values of $1-\alpha$. This is expected, as higher $1-\alpha$ typically results in non-singleton prediction sets, which produce a more conservative assignment of labels to cells, as discussed in Section III-B. Consequently, our planner generates longer paths to account for the most conservative labels. Similarly, since higher $1-\alpha$ often leads to non-singleton prediction sets, the proportion of the overall path spent on exploration is also higher for larger $1-\alpha$.
@@
Framework
$1-\alpha$
$S.R_{\text{map}}$
$S.R_{\text{mission}}$
Path
length
($m$)
Exploration
proportion
95 93.36 98.36 74.7 19
Ours 90 90.16 96.72 71.7 18.54
85 86.89 95.08 69.72 11.2
@@
UA
NA 10.17 65.57 48.91 NA
@@ V-D Comparative Experiments
Comparative Results: Table I shows the performance of both baselines using the evaluation metrics discussed in Section V-B. To ensure a fair comparison, both baselines are evaluated on the same $61$ test scenarios used for our proposed method. Notice that the UI planner outperforms the UA planner in terms of mission and mapping success rates, as it accounts for mapping uncertainty. However, our method outperforms the UI baseline because the UI prediction sets—unlike ours—are constructed heuristically and are not supported by any correctness guarantees.
@@
As seen in Table I, our approach leads to a significant improvement in both mapping and mission success rates. Our empirical evaluation shows that most baseline failures occur due to partial occlusions in the environment. An example is illustrated in Figure 4, where a human is partially occluded by trees. The cell associated with the human is often labeled as tree’ by the UA planner, leading to safety violations because the human’ label has stricter safety constraints than the tree’ label. In this case, by the time the UA planner detects the human, a safety constraint has already been violated. Similarly, the UI planner builds a singleton prediction set containing only the ‘tree’ label. In contrast, our approach constructs a prediction set that includes ‘tree’, ‘human’, and ‘free-space’, enabling the planner to follow a more conservative path and maintain safety. This results in higher mission and mapping success rates, though at the cost of longer paths, as shown in Table I.
Figure 4: Top view of the environment (left) and the corresponding 2D semantic map (right). The most likely label in ${\mathbf{m}}_{t}$ for the cell $j$ occupied by a human is ‘tree’. In contrast, the prediction set constructed by our method for that cell mitigates this mapping error by including ‘tree’, ‘human’, and ‘free-space’.
V-E Evaluation in Out-of-Distribution Scenarios
Finally, we evaluated our method on out-of-distribution (OOD) environments where the position and the number of trees are randomized (instead of fixed when $\mathcal{D}$} is used). We also changed the lighting conditions to simulate dawn and dusk. An example of an OOD environment is shown in Figure 5. For $\alpha=0.1$, the empirical mapping success rates dropped to $80.33\%$, but the mission completion rate was met at $93.44\%$. As expected, this shows that the guarantees may no longer hold in OOD settings; see Section VI.
Figure 5: Out-of-distribution environment with lighting that simulates dusk.
@@ VI Conclusions and Future Work
In this paper, we presented a novel semantic planner for unknown environments. Unlike related works, our method designs probabilistically correct paths in the presence of environmental uncertainty, stemming from imperfect perception, without any assumptions on sensor models or noise characteristics. The proposed method was supported both theoretically and experimentally. Future work will focus on: (i) designing active exploration algorithms that directly target reducing the size of prediction sets; (ii) accounting for OOD scenarios by leveraging robust CP methods; and (iii) directly quantifying sensor (rather than mapping) uncertainty.

### v2

@@ V-A Setting Up Experiments and the Proposed Planner
Distribution $\mathcal{D}$}: The distribution $\mathcal{D}$} is designed to generate environments as follows. All environments $\Omega$ have dimensions of $70\times 25$ m containing objects from five semantic classes $\mathcal{K}=\{\texttt{person},\texttt{car},\texttt{truck},\texttt{tree},\texttt{free}\}$. Each environment is populated with one car, one truck, one human placed at randomized positions and orientations, with the human biased toward being placed near trees, and three parallel rows of three trees in fixed locations. The class-based safety constraints are: $d_{\text{person}}=4m$, $d_{\text{car}}=1m$, $d_{\text{truck}}=2m$, and $d_{\text{tree}}=0.5m$. This choice enables testing our algorithm in cases of partial occlusions. An example environment is shown in Figure 3.
@@
Semantic Mapping: To set up the semantic mapping algorithm introduced in [ssmi_temp], we use a Detectron2 model [detectron2], trained on the COCO dataset [cocodataset], to semantically segment RGB images. The segmented outputs are fused over time to construct a 3D semantic octomap with resolution $r_{m}=1$ m. In our experiments, since we focus on a ground robot, we project this 3D map onto the X–Y plane, while considering height ranges relevant to the robot; this gives rise to a 2D semantic class–based grid map. In doing so, we assume that no semantic objects are stacked vertically.
@@
Exploitation-Exploration: The exploitation mode of Alg. 2 is implemented using a standard $\texttt{A}^{*}$ search algorithm on the 2D grid map [a-star]. For exploration, we design an algorithm that randomly selects a next state ${\mathbf{x}}_{t+1}$ near a voxel with a non-singleton prediction set subject to (6), aiming to decrease its prediction set size. As discussed in Section III-B, however, any other exploration algorithm can be substituted. [^3] A potential approach to handle cases in which the exploration algorithm fails to sufficiently reduce environmental uncertainty within a reasonable time is to find the largest $1-\alpha^{\prime}<1-\alpha$ for which a feasible path exists. This reduces the success rate to $1-\alpha^{\prime}$ but allows the proposed method to generate a best-effort path.
@@
Conformalization of Maps: Next, we discuss our CP implementation described in Section III-A. Recall that the NCS in (4) requires evaluating the maximum over the set $\mathcal{P}$}, which contains an infinite number of paths. Since this is impractical, following [PwC], we approximate the NCS using a finite subset $\hat{\mathcal{P}}\subset{\mathcal{P}}$ consisting of 10 paths between start and goal locations that are randomly selected in the calibration environment $\Omega_{i}\sim{\mathcal{D}}$. At test time, given $\Omega_{\text{test}}\sim{\mathcal{D}}$, we randomly select the start and goal positions for the task $\psi$. Also, to avoid generating a new calibration dataset for every $\Omega_{\text{test}}$, we apply CP as described in Remark IV.4, with $\delta=0.1$. Our calibration dataset consists of $50$ environments.
@@
Evaluation Metrics: We consider four metrics. (i) Mapping success rate, defined as the percentage of cases, where the following binary variable is $1$: $S.R_{\text{map}}:=\mathds{1}(C_{0:H}),$ where $C_{0:H}=\bigwedge_{\forall j\in{\mathcal{J}}_{t},t\in[0:H]}\left(k_{j}^{\text{gt}}\in{\mathcal{C}}_{j}({\mathbf{m}}_{t})\right)$ and $k_{j}^{\text{gt}}$ is the ground truth label of cell $j$. In words, for a given test environment $\Omega_{\text{test}}$, $S.R_{\text{map}}=1$ if all constructed prediction sets ${\mathcal{C}}_{j}({\mathbf{m}}_{t})$ contain the true labels $k_{j}^{\text{gt}}$ for all $j\in{\mathcal{J}}_{t}$ and $t\in[0:H]$. Due to Proposition IV.1, the average mapping success rate of Alg. 2 across test environments drawn from $\mathcal{D}$} is expected to be at least $1-\alpha$ even in the presence of a sensor with limited FoV. (ii) Mission success rate, defined as the percentage of cases, where the binary variable $S.R_{\text{mission}}:=\mathds{1}\left({\mathbf{x}}_{0:H}\models\psi\right)$ is $1$. In words, given a test environment $\Omega_{\text{test}}$, and task $\psi_{\text{test}}$, $S.R_{\text{mission}}=1$ if the robot path ${\mathbf{x}}_{0:H}$ satisfies the safety constraints in (9) and ${\mathbf{x}}_{H}\in{\mathcal{X}}_{g}$. Note that if the sensor was omnidirectional, the mission success rate of Alg. 2 should have been to be at least $1-\alpha$ due to Theorem IV.3; see our discussion in Section V-B. (iii) Distance traveled by the robot during deployment, which captures efficiency of the generated paths. (iv) Proportion of the overall path used for exploration, denoted as $p_{\text{expl}}$, quantifying how much of the trajectory was dedicated to reducing uncertainty in the environment. We report the average of these four metrics over 61 test environments. We also report the standard deviation of the binary variables $S.R_{\text{map}}$ and $S.R_{\text{mission}}$.
V-B Empirical Evaluation of the Proposed Planner
@@
We evaluate our proposed method for three user-defined success rates: $1-\alpha\in\{95\%,90\%,85\%\}$ over $61$ environments drawn from $\mathcal{D}$}. The empirical results are reported in Table I. Observe that the empirical mapping success rates validate Prop. IV.1. It can be seen that for $1-\alpha=95\%$, the empirical mapping success rate is $93.36\%<95\%$. Such small deviations are due to the finiteness of the calibration dataset, the number of paths that are considered for calibration, and the number of validation environments; see [CPbase] for more information on the influence of finiteness on the empirical success rates. We also observe that the empirical mission success rates validate Theorem IV.3. We note that we did not observe any mission failures due to limited FoV.  [^4] A heuristic approach to prevent mission failures due to limited FoV is to have the robot spin at time $t$ to collect measurements, or to restrict motions into unknown areas lying within its blind spots. The path length increases with higher values of $1-\alpha$. This is expected, as higher $1-\alpha$ typically results in non-singleton prediction sets, which produce a more conservative assignment of labels to cells, as discussed in Section III-B. Consequently, our planner generates longer paths to account for the most conservative labels. Similarly, since higher $1-\alpha$ often leads to non-singleton prediction sets, the proportion of the overall path spent on exploration is also higher.
@@
Planner $1-\alpha$ (%) $S.R_{\text{map}}$ (%) $S.R_{\text{mission}}$ (%) Path length ($m$) $p_{\text{expl}}$ (%)
95 93.36 $\pm$ 3.19 98.36 $\pm$ 1.63 74.7 19
Ours 90 90.16 $\pm$ 3.81 96.72 $\pm$ 2.28 71.7 18.54
85 86.89 $\pm$ 4.32 95.08 $\pm$ 2.77 69.72 11.2
@@
UA NA 10.17 65.57 48.91 NA
[…]
$1-\alpha$ (%) $n$ $S.R_{\text{map}}$ (%) $S.R_{\text{mission}}$ (%) Path length ($m$) $p_{\text{expl}}$ (%)
1 90.33 $\pm$ 6.04 96.56 $\pm$ 2.34 63.727 5.59
5 87.66 $\pm$ 4.12 89.22 $\pm$ 3.94 75.206 13.59
90 10 90.16 $\pm$ 3.81 96.72 $\pm$ 2.28 71.7 18.54
20 90 $\pm$ 3.79 100 $\pm$ 0 76.7338 18.81
Figure 4: Top view of the environment (left) and the corresponding map ${\mathbf{m}}_{t}$ (right). The most likely label in ${\mathbf{m}}_{t}$ for cell $j$, occupied by a human, is ‘tree’. In contrast, the prediction set constructed by our method for that cell mitigates this mapping error by including ‘tree’, ‘human’, and ‘free-space’.
@@ V-D Comparative Experiments
Comparative Results: Table I shows the performance of both baselines using the evaluation metrics discussed in Section V-A. To ensure a fair comparison, both baselines are evaluated on the same $61$ test environments, and their corresponding reach-avoid tasks used for our proposed method. Notice that the UI planner outperforms the UA planner in terms of mission and mapping success rates, as it accounts for mapping uncertainty. However, our method outperforms the UI baseline because the UI prediction sets—unlike ours—are constructed heuristically.
@@
As seen in Table I, our approach leads to a significant improvement in mapping and mission success rates. Our empirical evaluation shows that most baseline failures occur due to partial occlusions in the environment. An example is illustrated in Figure 4, where a human is partially occluded by trees. The cell associated with the human is often labeled as tree’ by the UA planner, leading to safety violations because the human’ label has stricter safety constraints than the tree’ label. In this case, by the time the UA planner detects the human, a safety constraint has already been violated. Similarly, the UI planner builds a singleton prediction set containing only the ‘tree’ label. In contrast, our approach constructs a prediction set that includes ‘tree’, ‘human’, and ‘free-space’, enabling the planner to follow a conservative path and maintain safety. This results in higher mission and mapping success rates, though at the cost of longer paths, as shown in Table I.
Figure 5: Hardware demonstration of the proposed method.
V-E Hardware Experiments
In this section, we demonstrate robot behaviors with respect to various values for $1-\alpha$ in real-world settings. Specifically, we consider a Clearpath Jackal equipped with an Intel RealSense D455 camera and a Detectron2 segmentation model [detectron2]. The robot operates in environments with humans, motorcycles, cars, and tree (see Fig. 5). The class-based safety constraints are: $d_{\text{person}}=1.5m$, $d_{\text{car}}=1.25m$, $d_{\text{motorcycle}}=1m$, and $d_{\text{tree}}=1m$. Similar to the simulation results, we observed an increasing trend in path length as $1-\alpha$ increases. Specifically, the total traveled distance for $1-\alpha=90\%,95\%,\text{and }99\%$ was $24.38$ m, $24.92$ m, and $25.78$ m, respectively. We note that our method was applied using calibration data described in Section V-A, which induces a distribution shift. Evaluation of our approach in OOD settings in terms of mapping and mission success rates is provided in Section V-F.
V-F Evaluation in Out-of-Distribution (OOD) Scenarios
Finally, we evaluated our method in OOD environments. For calibration, we use the same $50$ calibration environments sampled from $\mathcal{D}$} mentioned in Section V-A, with $|\hat{{\mathcal{P}}}|=10$. In what follows, we set $1-\alpha=90\%$. We consider the case where the test environments are sampled from a distribution $\mathcal{D}$$^{\prime}$ where, unlike $\mathcal{D}$}, the trees vary in number from $1$ to $10$, and are placed randomly. In addition, the environments also simulated the lighting conditions of dusk and dawn instead of midday used in $\mathcal{D}$}. We observed a drop in mapping success rates for $1-\alpha=90\%$ from $90.16\%\pm 3.81\%$ to $80.33\%\pm 5.09\%$. The mission success rates also dropped from $96.72\%\pm 2.28\%$ to $93.44\%\pm 3.17\%$, albeit still being greater than $1-\alpha$. We empirically observed that changes in the number and spatial distribution of the trees have a more significant impact than the lightning conditions on the performance of the algorithm.
@@ VI Conclusions and Future Work
In this paper, we presented a novel semantic planner for unknown environments. Unlike related works, our method designs probabilistically correct paths in the presence of environmental uncertainty, stemming from imperfect perception, without any assumptions on sensor models or noise characteristics. Future work will focus on: (i) designing active exploration algorithms that directly target reducing the size of prediction sets; (ii) accounting for OOD scenarios by leveraging robust CP methods; and (iii) directly quantifying sensor (rather than mapping) uncertainty.